solve(D^2+2DD'+D'^2)z = e^x+2y + sinh(x+y)
1
Expert's answer
2020-07-07T21:13:03-0400

"\left(D^{2}+2 D D^{\prime}+D^{\prime 2}\right) z=e^{(x+2 y)}+\sinh(x+y)"


The auxiliary equation is "m^{2}+2m + 1 = 0"

"\begin{array}{l}
(m+1)^{2}=0 \\
\Rightarrow m_{1}=-1, m_{2}=-1 \\
\therefore \text{The complementary function}\\
C.F = f_{1}(y - x) + x f_{2}(y - x) \\
\end{array}"


Let, the particular integral be


"P.I = P.I_{1}+P.I_{2}\\
\begin{aligned}
P.I_{1}&=\frac{1}{D^{2}+2 DD'+D'^{2}} \cdot e^{(x+2 y)} \quad\left(D = 1, D' = 2\right)\\
P.I_{1}&=\frac{1}{9} e^{(x+2y)}
\end{aligned}"

"\begin{aligned}
P.I_{2}&=\frac{1}{D^{2}+2 D D^{\prime}+D^{2}} \cdot \sinh (x+y) \\
&=\frac{1}{D^{2}+2 D D^{\prime}+D^{2}}\left[\frac{e^{(x+y)}-e^{-(x+y)}}{2}\right]\\
&=\frac{1}{2}\left[\frac{1}{D^{2}+2 D D^{\prime}+D^{\prime 2}} \cdot e^{(x+y)}-\frac{1}{D^{2}+2 D D^{\prime}+D^{\prime 2}} \cdot e^{-(x+y)}\right]\\
& \qquad\qquad \{D = 1, D' = 1\} \qquad\qquad\qquad \{D =-1, D'=-1\}\\
&=\frac{1}{2}\left[\frac{e^{x+y}}{4}-\frac{e^{-(x+y)}}{4}\right]\\
P.I_{2}&=\frac{1}{8}\left[e^{(x+y)}-e^{-(x+y)}\right]=\frac{1}{4} \sinh (x+y)
\end{aligned}"


The complete solution is,

"z=f_{1}(y-x)+x f_{2}(y-x)+\dfrac{e^{x+2 y}}{9}+\dfrac{1}{4} \sinh (x+y)"